OpenStudy (beatboxingfetus):

Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (beatboxingfetus):

Tell me if you're not able to see the picture.

OpenStudy (beatboxingfetus):

@Ashes.boom ( sorry for asking you so much q's)

OpenStudy (ashes.boom):

its okay :)

OpenStudy (anonymous):

couldn't see the pic

OpenStudy (ashes.boom):

i couldnt se it either

OpenStudy (beatboxingfetus):

there you go

OpenStudy (beatboxingfetus):

c:

OpenStudy (ashes.boom):

okay start with y intercept. What point is the y intercept?

OpenStudy (beatboxingfetus):

3

OpenStudy (beatboxingfetus):

oh waiton the graph? xD

OpenStudy (ashes.boom):

technically 300 yes

OpenStudy (ashes.boom):

haha yea

OpenStudy (beatboxingfetus):

so they y intercept is 300?

OpenStudy (ashes.boom):

yes

OpenStudy (ashes.boom):

so now you have y = mx + 300

OpenStudy (ashes.boom):

remmebr how to find the slope

OpenStudy (beatboxingfetus):

the slope is 150?

OpenStudy (ashes.boom):

slope is a fraction becuase the formula is y2 - y1 / x2 - x1 so use the y intercept as one point and find another point to use for the second one

OpenStudy (beatboxingfetus):

hmm??

OpenStudy (anonymous):

actually its x2-x1 / y2 - y1 , am i right ?

OpenStudy (ashes.boom):

the y goes first i think

OpenStudy (ashes.boom):

okay so the y intercept is at point 0, 300

OpenStudy (ashes.boom):

the other point i see is 150, 200

OpenStudy (ashes.boom):

200-300 over 150 - 0

OpenStudy (ashes.boom):

100/150

OpenStudy (ashes.boom):

100/150 = 10/15 = 2/3

OpenStudy (ashes.boom):

the line is going down so its subtraction -2/3

OpenStudy (ashes.boom):

y = -2/3x + 300

OpenStudy (beatboxingfetus):

man i really dislike math. Thank you so much for helping me with my assignment, i would follow you ashes but idk how bc open study changed stuff and it's not like popping your name up when i hoover my mouse over it.

OpenStudy (beatboxingfetus):

sorry my internet went out!